OpenStudy (anonymous):

PLEASE HELP!!! MEDALS AND I'LL FOLLOW YOU!!! Julie bought a drink for $2.99, two bags of chips for $1.29 each, and some gasoline at a convenience store. The price of gasoline was $2.39 per gallon. If x represents the number of gallons of gasoline she purchased, which of the following equations can be used to find how much Julie spent at the convenience store? Answers down below!!

OpenStudy (anonymous):

A. y = 2.39x + 4.28 B. y = 2.39x + 5.57 C. y = 4.28x + 2.39 D. y = 5.57x + 2.39

OpenStudy (anonymous):

@jesusfreak565 @Yuba